Find the domain of the function:
g(x) = 6/8-7x

Answer by midwood_trail(310)

Find the domain of the function:
g(x) = 6/(8-7x)
The domain of this function will be all real numbers except the number that will produce division by zero. Division by zero is not allowed or does not exist. In your math book you'll see the word UNDEFINED. This word means DOES NOT EXIST.
If we replace x with the improper fraction 8/7, we create division by zero in the denominator of this fraction which is not allowed.
This is what I mean:
Denominator: 8 - 7x
Let x = 8/7
Then: 8 - 7(8/7) = 8 - 8 = 0 in the denominator.
Here is the domain:
The domain is ALL REAL NUMBERS EXCEPT that x CANNOT = 8/7.
Did you follow?